OpenStudy (b2st):

A car starts from rest and travels for 5.2 s with a uniform acceleration of +1.9 m/s2. The driver then applies the brakes, causing a uniform acceleration of -1.6 m/s2. The breaks are applied for 1.90 s. (b) How far has the car gone from its start? This is the last question I need to do, and I'm like super stuck because I keep getting the answer wrong.

OpenStudy (b2st):

I understand that I do 2 separate problems and add them up together later, but I probably did something wrong. Perhaps plugging in the wrong numbers in?

OpenStudy (loser66):

what do you have?

OpenStudy (b2st):

Okay I used this formula : x = v1(t) + 1/2(a)(t)^2

OpenStudy (b2st):

correct right? Uhm, I used v1 = 0 (not sure about this) & lpug the rest of the numbers in.

OpenStudy (loser66):

ok,

OpenStudy (b2st):

Is that correct?

OpenStudy (loser66):

number? \(x_1=\)

OpenStudy (b2st):

??

OpenStudy (b2st):

Oh the first x1 term I got 16.128

OpenStudy (loser66):

how? \(x_1 = \dfrac {1}{2}*(1.9)*(5.2)^2=25.688\)

OpenStudy (b2st):

oh opps. I was doing the practice problem. Haha my bad.

OpenStudy (b2st):

But yeah I got that answer from my previous one.

OpenStudy (loser66):

ok,next?

OpenStudy (b2st):

-1.52

OpenStudy (b2st):

opps

OpenStudy (loser66):

you have to find \(v_2\)

OpenStudy (b2st):

-2.8

OpenStudy (b2st):

v2? as velocity final?

OpenStudy (b2st):

Because for part (a) for the question it was asking for the velocity final and I got 6.84 m/s (which is correct)

OpenStudy (loser66):

you start at rest and get \(V_2\) then, from \(V_2\) you apply braking to stop.

OpenStudy (loser66):

it's not final, final is \(V_3 =0\) the time you have full stop

OpenStudy (b2st):

Ohh, I see what I did wrong. So for the second for 1.9s & -1.6 m/s^2, I use 6.84? or the v2 for the second part.

OpenStudy (loser66):

how can you get 6.84?

OpenStudy (b2st):

This is a from a previous question. Asking for the final velocity for total. So I got 6.84 m/s

OpenStudy (b2st):

(a) How fast is the car going at the end of the braking period? -> answer 6.84 m/s

OpenStudy (b2st):

I use a different thing, I don't use v1 or v2 in my class. Instead I use Vf and VI

OpenStudy (loser66):

I don't. I apply \[V_2 = V_1 + at \] where \(V_1 =0 ~~at~~rest\) \[V_2 = 1.9* 5.2 = 9.88 m/s

OpenStudy (b2st):

Okay, so I used that and plug for v1 for the second equation??

OpenStudy (loser66):

nope, for this part I apply \[V_3^2 = V_2^2 +2as \\s = \frac{V_3^2 -V_2^2}{2a}\]this a is -1.6

OpenStudy (b2st):

@ ^ @ the numbers are confusing. I have trouble organizing numbers. lol Can you possibly do vf or vi? It's easier for me know where to go.

OpenStudy (loser66):

and I got \(x_2 = 30.50 m\)

OpenStudy (loser66):